How deep should be ocean to enhance physiologically harmful infrasonic waves of 6 Hz frequency? Take 1.47 km/s for the speed of sound in water and find the minimum and next to minimum depths of the ocean.

The minimum depth, h1 =  

The next to minimum depth, h2 =

What is the length of the string if the speed of wave is 255 m/s?

The length of the string, L =

The frequency of a radio station is 88 MHz. Find the period and wavelength of the radio waves emitted by this station in empty space. The speed of electromagnetic waves in vacuum is c=3⋅108c=3⋅108 m/s.

The period, T =

(Answer All Pleaseeee)
Answer #1

3.

f=88*10^6 Hz

Use formula for period

T=1/f=1/(88*10^6 Hz)

T=1.14*10^-8 s

Use formula for wavelength,

v=fλ

λ=v/f

λ=(3*10^8m/s)/(88*10^6 Hz)

λ=3.41 m